# Rate-reliability functions

### v2

@@ III-C Improved bounds for subsets of error exponents
In Corollaries 5 and 8 we have lower- and upper-bounded the rate-reliability function in terms of the lower and upper Minkowski dimensions, respectively. These bounds are universal in that they hold for all error exponents $E$ smaller than some threshold $E_{0}=E_{0}(\eta)$ that decreases with the parameter $\eta>0$. We might however be interested only in certain (infinitely many) sufficiently small values of $E$. To capture this, we consider subsets $\mathcal{E}\subset(0;+\infty)$ that have 00 as an accumulation point (meaning that $\mathcal{E}$ contains sequences converging to 00).
@@
Substituting the above expression into Eq. (9), we obtain an improved achievability bound in Corollary 5, but now with the upper Minkowski dimension and under the added condition that the error exponents $E(n)\in\mathcal{E}_{g}$.
@@
$\lim_{\mathcal{E}_{b}\ni E\rightarrow 0}\frac{\log\left[\Gamma_{\frac{1}{2}\sqrt{1-e^{-E/2}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]}{-\log\left(\frac{1}{2}\sqrt{1-e^{-E/2}}\right)}=\underline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$
@@
where we have followed the last steps in the proof of Corollary 8, using again the expansion in Eq. (19). Plugging the above expression into Eq. (15), we obtain an improved upper bound on the rate-reliability function in Corollary 8, but now with the lower Minkowski dimension and under the added condition that the error exponents $E(n)\in\mathcal{E}_{b}$.
IV Discussion: constant vs vanishing exponents, dimension zero, and quantum channels
In this section, we demonstrate how the new reliability framework developed in this study aligns with the established results in [CDBW:DI_classical]. We begin by deriving bounds for the rate-reliability functions in the asymptotic regime ($n\rightarrow\infty$), showing how these bounds allow us to recover the main capacity results from [CDBW:DI_classical]. Next, we consider the special case where the Minkowski dimension of the output probability set is zero – a scenario where Corollaries 5 and 8 do not fully capture the underlying behavior meaningfully. We illustrate an approach to handle such cases with a straightforward example. Finally, we extend our results to both classical-quantum and (fully) quantum channels, under the restriction that encoding is limited to product states. Together, these sections encompass the contents of [CDBW:DI_classical] through reliability arguments, offering a new perspective on deterministic identification.
IV-A From linear to superlinear DI rates
The results from the preceding section tell us that the superlinear rates of DI codes discussed in the introduction are lost when imposing exponentially vanishing errors. Indeed, when $E_{1},\,E_{2}>0$, then Corollary 8 shows for every $\eta>0$, letting $E=\min\{E_{1},E_{2},E_{0}(\eta)\}$, that
@@ IV-A From linear to superlinear DI rates
This is because we can take the limit in Eq. (17) for every parameter $\eta>0$, as long as $E$ is below the cutoff exponent $E_{0}(\eta)$; at the same time, the DI rate can only increase if we lower $E$.
Conversely, and similarly, for every $\eta>0$ and $10<t<$ 1, and $E>0$ sufficiently small, Corollary 5 implies for all $E_{1},\,E_{2}\leq E$ that
@@
which is positive for sufficiently small $E>0$, and indeed diverges to infinity as $E\rightarrow 0$.
On the other hand, the general superlinear scaling of DI codes featured in [CDBW:DI_classical] does reappear when we consider (slowly) vanishing error exponents $\lim_{n\rightarrow\infty}E_{1}(n)=\lim_{n\rightarrow\infty}E_{2}(n)=0$. Evidently, we still need small errors $\lambda_{1},\,\lambda_{2}\ll 1$ as $n\rightarrow\infty$ to have a good code, so the $nE_{i}(n)$ must remain bounded away from 00; more precisely, we need $E_{1}(n),\,E_{2}(n)\geq\omega(1/n)$ for the coding part and weak converse, and $E_{1}(n),\,E_{2}(n)\geq\Omega(1/n)$ for the strong converse. By direct substitution of these two “settings” into the results above we can recover the results from [CDBW:DI_classical]. We point out that those results are bounds on the superlinear capacity (the maximum asymptotic rate), which we can recover as follows in the present notation:
@@
where the inner maximum is over all DI codes with errors of first and second kind bounded as $\lambda_{i}\leq 2^{-nE_{i}(n)}$, and the outer supremum is over functional dependencies of the exponents on $n$ going to 00 sufficiently slowly: $\omega(1/n)\leq E_{i}(n)\leq o(1)$.
Theorem 9 (Cf. [CDBW:DI_classical, Thm. 5.7]).
The superlinear DI capacity $\dot{C}_{\text{DI}}(W)$ for the channel $W:\mathcal{X}\rightarrow\mathcal{Y}$ is bounded in terms of the lower Minkowski dimension of the set $\sqrt{\!\widetilde{\mathcal{X}}}$ as follows (for slowly vanishing error exponents $\lim_{n\rightarrow\infty}E_{i}(n)=0$):
$\frac{1}{4}\underline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\leq\dot{C}_{\text{DI}}(W)\leq\frac{1}{2}\underline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$
@@
We start with the coding part taking $E_{1}(n)=E_{2}(n)=\frac{\log n}{n}$ in Eq. (14) (Corollary 5):
$\begin{split}R(n)&\geq\frac{1-t}{4}\left[\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)-\eta\right]\log\frac{ct^{2}n}{6\log n}\\ &\phantom{\geq.}-H(t,1\!-\!t)-O\!\left(\frac{\log n}{n}\right).\end{split}$
@@
$\dot{C}_{\text{DI}}(W)\geq\liminf_{n\rightarrow\infty}\frac{1}{\log n}R(n)\geq\frac{1}{4}\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right),$
@@
Similarly, for the upper bound we need to take $\mathcal{E}_{b}\ni E_{1}(n)=E_{2}(n)\geq\frac{C}{n}$ in Eq. (21), the (pessimistic) upper bound for bad sequences of errors. Choosing also $\eta=\eta(n)$ going slowly enough towards 00 for increasing $n$, such that the error exponents, which are upper bounded by the threshold $E_{0}\!\left(\eta(n)\right)$, can also go to 00 slowly, we find
$\begin{split}\dot{C}_{\text{DI}}(W)&\leq\liminf_{n\rightarrow\infty}\frac{1}{2\log n}\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\log\frac{8n}{C}+O\left(\frac{1}{n}\right)\\ &=\frac{1}{2}\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right),\end{split}$
@@
We refer the reader to [CDBW:DI_classical, Sect. 7.1] for an extensive discussion on the origins, comparison, and interpretation of the optimistic and pessimistic capacities, or else [Ahlswede2006] which inspired it. We can recover the general superexponential optimistic capacity bounds in [CDBW:DI_classical] with similar methods as in the previous theorem.
Theorem 10 (Cf. [CDBW:DI_classical, Thm. 5.9]).
The optimistic superlinear DI capacity $\dot{C}_{\text{DI}}^{\text{opt}}(W)$ of a channel $W:\mathcal{X}\rightarrow\mathcal{Y}$ is bounded in terms of the upper Minkowski dimension of $\sqrt{\!\widetilde{\mathcal{X}}}$ as follows (for slowly vanishing error exponents $\lim_{n\rightarrow\infty}E_{i}(n)=0$):
$\frac{1}{4}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\leq\dot{C}_{\text{DI}}^{\text{opt}}(W)\leq\frac{1}{2}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$
@@
We follow the same steps as in the previous proof. For the direct part we need to take $\mathcal{E}_{g}\ni E_{1}(n),\,E_{2}(n)\geq\frac{C}{n}$ in the (optimistic) lower bound for good sequences of error exponents [Eq. (20)], and similarly choose $t=t(n)$ and $\eta=\eta(n)$ going to 00 simultaneously and sufficiently slowly as $n\rightarrow\infty$. Then,
$\begin{split}\dot{C}_{\text{DI}}^{\text{opt}}(W)&\geq\limsup_{n\rightarrow\infty}\frac{1}{4\log n}\left[\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)-\eta(n)\right]\log\frac{ct^{2}n}{6C}\\ &=\frac{1}{4}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).\end{split}$
For the converse we take $E_{1}(n)=E_{2}(n)=\frac{\log n}{n}$ in Eq. (17) and $\eta=\eta(n)$ going to zero slowly enough, and find
$\begin{split}\dot{C}_{\text{DI}}^{\text{opt}}(W)&\leq\limsup_{n\rightarrow\infty}\frac{1}{2\log n}\left[\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)+\eta(n)\right]\log\frac{8n}{\log n}\\ &\phantom{\leq.}+O\left(\frac{1}{n}\right)=\frac{1}{2}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right),\end{split}$
@@
We have just seen that we can recover the capacity bounds in [CDBW:DI_classical] from the rate-reliability bounds derived above. However, we have to impose a slow decrease of the errors in block length. The convergence of the rate bounds towards the capacity is also very slow, as shown in Figure 2.
@@ IV-B Zero-dimensional output probability sets
In Corollaries 5 and 8, and in Section III-C, we have demonstrated how to bound the rate $R(n)$ of a code with small error exponents using the lower and upper Minkowski dimensions featured in the main pessimistic and optimistic capacity results of [CDBW:DI_classical]. This could cause some confusion on how to approach channels for which the output probability set (equivalently its square-root set) has Minkowski dimension $d_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)=0$, particularly in the regime of small error exponents, where the aforementioned corollaries apply. Indeed, in such a case, the lower bound given in Corollary 5 becomes trivial, and the upper bound only tells us that $R(n)\leq o\left(\log\frac{1}{E(n)}\right)$ for small $E(n)\leq E_{1}(n),\,E_{2}(n)$. However, we can obtain much more meaningful information if we go back to Theorems 4 and 7 (for the lower and upper bounds respectively) and impose the regime of small errors only at an appropriate point.
@@
$B(y|x)=B_{x}(y)=xy+(1-x)(1-y)=\begin{cases}x&\text{for }y=1,\\ 1-x&\text{for }y=0.\end{cases}$
Example 11 (Cf. [CDBW:DI_classical, Subsect. 5.4]).
For a real number $a>1$, let $\mathcal{X}_{a}:=\{0\}\cup\{a^{-k}:0\leq k\in\mathds{N}_{0}\}$ be a set of numbers in the interval $[0,1]$. Then, the maximum rate $R(n)$ of a DI-code for $B|_{\mathcal{X}_{a}}$, the Bernoulli channel restricted to inputs from $\mathcal{X}_{a}$, for small and equal error exponents $E(n):=E_{1}(n)=E_{2}(n)>0$, is bounded as follows:
$\log\log\frac{1}{E(n)}-O(1)\leq\max R(n)\leq\log\log\frac{1}{E(n)}+O(1).$
@@
As the set of all probability distributions with binary output $\mathcal{P}(\{0,1\})$ (with total variation distance) is isometric to the interval $[0;1]$ (with the usual metric on real numbers), the covering (or packing) of the output probability set of a channel with binary output (like $B|_{\mathcal{X}_{a}}$) corresponds to the covering (or packing) of its input, which is $\mathcal{X}_{a}$. It is not hard to see that
@@
To prove the lower bound, we start from Theorem 4, we use the covering/packing relation $\Pi_{\epsilon}(F)\geq\Gamma_{2\epsilon}(F)$, and apply the lower bound in Eq. (24) (noting that in this example $\sqrt{\mathcal{X}_{a}}=\mathcal{X}_{\sqrt{a}}$). After some basic algebra we get that there exists a DI code with
$\begin{split}R(n)&\geq(1-t)\log\log_{a}\frac{t\sqrt{c}(\sqrt{a}-1)^{2}}{36\sqrt{6E(n)}}\\ &\phantom{\geq.}-H(t,1-t)-O\left(\frac{\log n}{n}\right).\end{split}$
To optimise the right hand side, we choose $t=\sqrt[4]{E(n)}$, such that
@@
For the upper bound we start from Theorem 7 and use the right-hand inequality in Eq. (24) and the property $\sqrt{\mathcal{X}_{a}}=\mathcal{X}_{\sqrt{a}}$:
$\begin{split}R(n)&\leq\log\left[\Gamma_{\frac{1}{2}{\sqrt{1-e^{-E(n)/2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]\\ &\leq\log\log_{\sqrt{a}}\left(\frac{a}{\sqrt{1-e^{-E(n)/2}}}\right).\end{split}$
@@
$\begin{split}R(n)&\leq\log\log_{a}\frac{2a^{2}}{E(n)}+O\left(\frac{E(n)}{-\log E(n)}\right)\\ &=\log\log\frac{1}{E(n)}+O(1).\end{split}$
Considering the bounds above and the arguments in Section IV-A, where we took $E(n)=\omega(1/n)$ for the lower bound and $E(n)=\Omega(1/n)$ for the upper, we conclude that the DI capacity in this case ought be defined at $n\log\log n$ scale (cf. [CDBW:DI_classical, Subsect. 5.4]), and that it evaluates to
@@
The linear capacity results for the DMC from [AC:DI, SPBD:DI_power] are recovered from these two bounds in the limit of vanishing exponents: $C_{\text{DI}}(W)=\log|W(\mathcal{X})|$.
@@
For the direct part [Eq. (25)] we can reuse the proof of Theorem 4 with slight modifications. There, we started by defining a packing on $\sqrt{\!\mathcal{X}}$ of radius $\beta$, and worked after with the discrete set of $|\mathcal{X}_{0}|$ packing elements. Since now we already have a finite set of output probability distributions $\widetilde{\mathcal{X}}=W(\mathcal{X})$, we let
$2\beta:=\min_{x\neq x^{\prime}\in\mathcal{X}}\left|\sqrt{W_{x}}-\sqrt{W_{x^{\prime}}}\right|_{2},$
the minimum Euclidean distance between square roots of output probability distributions, and the rest of the proof is recycled. We will create a maximum-size code in block length $n$ with minimum Hamming distance between different code words $d_{H}(x^{n},x^{\prime n})\geq tn$ and partition the code into regions of similar entropy. Following the arguments, we can lower-bound the rate as [cf. Eq. (13)]:
$R(n)\geq(1-t)\log|W(\mathcal{X})|-H(t,1-t)-O\left(\frac{\log n}{n}\right),$
@@
$R(n)\geq\log|W(\mathcal{X})|-O\left(-\sqrt{E(n)}\log E(n)\right)-O\left(\frac{\log n}{n}\right).$
For the converse [Eq. (26)] we start by noticing that Eq. (16) translates to a minimum Hamming distance between code words. Namely, letting $\frac{1}{\alpha}:=\displaystyle\max_{x\neq x^{\prime}\in\mathcal{X}}F(W_{x},W_{x^{\prime}})<1$ be the maximum fidelity between different output probability distributions, we have for two distinct code words $u_{j}=x^{n}$ and $u_{k}={x^{\prime}}^{n}$:
$\frac{1}{\alpha^{d_{H}(u_{j},u_{k})}}\leq 4e^{-nE(n)},$
@@
$R(n)\leq\log|W(\mathcal{X})|-O\left(-E(n)\log E(n)\right).$
The only thing left to show is that in the asymptotic limit, the rate bounds above yield the linear capacity results from [SPBD:DI_power]. Repeating the argument in Section IV-A we take $E(n)\geq\Omega(1/n)$ for the upper bound and $E(n)\geq\omega(1/n)$ for the lower, which translates into:
$\begin{split}\max R(n)&\geq\log|W(\mathcal{X})|-O\left(\frac{1}{\sqrt{n}}\log n\right)\\ \max R(n)&\leq\log|W(\mathcal{X})|-O\left(\frac{1}{n}\log n\right).\end{split}$

### v4

@@ III-C Improved bounds for subsets of error exponents
In Corollaries 6 and 8 we have lower- and upper-bounded the rate-reliability function in terms of the lower and upper Minkowski dimensions, respectively. These bounds are universal in the sense that they hold for all error exponents $E$ smaller than some threshold $E_{0}=E_{0}(\eta)$ that decreases with the arbitrary parameter $\eta>0$. We might however be interested only in certain (infinitely many) sufficiently small values of $E$. To capture this, we consider subsets $\mathcal{E}\subset(0;+\infty)$ that have 0 as an accumulation point (meaning that $\mathcal{E}$ contains sequences converging to 0).
@@
Substituting the above expression into Equation (17), we obtain an improved achievability bound in Corollary 6, but now with the upper Minkowski dimension and under the added condition that the error exponents $E(n)\in\mathcal{E}_{g}$.
@@
$\lim_{\mathcal{E}_{b}\ni E\rightarrow 0}\frac{\log\left[\Gamma_{\frac{1}{2}\sqrt{1-e^{-E/2}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]}{-\log\left(\frac{1}{2}\sqrt{1-e^{-E/2}}\right)}=\underline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$ (40)
@@
where we have followed the last steps in the proof of Corollary 8, using again the expansion in Equation (37). Plugging the above expression into Equation (27), we obtain an improved upper bound on the rate-reliability function in Corollary 8, but now with the lower Minkowski dimension and under the added condition that the error exponents $E(n)\in\mathcal{E}_{b}$. These improved bounds for good and bad subsets of error exponents will be realted to the optimistic and pessimistic capacity approaches respectively in the next section, specifically in IV-A.
IV Discussion: constant vs vanishing exponents, dimension zero, and large exponent regime
In this section, we demonstrate how the new reliability framework developed in this study can yield the established results in [17]. We begin by deriving bounds for the rate-reliability functions in the asymptotic regime ($n\rightarrow\infty$), showing how these bounds allow us to recover the main established capacity results under some particular error conditions. Next, we consider the special case where the Minkowski dimension of the output probability set is zero—a scenario where Corollaries 6 and 8 do not fully capture the underlying behaviour meaningfully. We illustrate an approach to handle such cases with a couple of straightforward examples. Finally, in IV-C, we discuss the regime of large error exponents which, as we have already hinted in 5, is not bounded meaningfully through the presented Theorems.
IV-A From linear to linearithmic DI rates
The results from the preceding section tell us that the superlinear rates which, as discussed in the introduction, are a defining characteristic of general DI codes, are lost when imposing exponentially vanishing errors. Indeed, in the asymptotic limit $n\rightarrow\infty$, let $E=\min\{E_{1},E_{2},E_{0}(\eta)\}>0$, then Corollary 8 shows that for every $\eta>0$:
@@ IV-A From linear to linearithmic DI rates
This is because we can take the limit in Equation (35) for every parameter $\eta>0$, as long as $E$ is below the cutoff exponent $E_{0}(\eta)$.
[…]
@@
which is positive for sufficiently small $E>0$. Indeed, Equations (42) and (43) tell us that the linear-scale rate for DI is a bounded positive number under exponentially vanishing errors (the condition under which the previous Section III is developed) with $E>0$.
At the same time, however, we notice that the DI rates increase when we lower $E$. Actually, both Equations (42) and (43) diverge to infinity as $E\rightarrow 0$! Meaning that the rate scaling is faster than linear in this regime. As a matter of fact, we show next that the general linearithmic scaling of DI codes featured in [17] does reappear when we consider (slowly) vanishing error exponents.
Let $\lim_{n\rightarrow\infty}E_{1}(n)=\lim_{n\rightarrow\infty}E_{2}(n)=0$. Evidently, we still need small errors $\lambda_{1},\,\lambda_{2}\ll 1$ as $n\rightarrow\infty$ to have a good code, so the $nE_{1,2}(n)$ must remain bounded away from 0; more precisely, we need $E_{1}(n),\,E_{2}(n)\geq\omega(1/n)$ for the coding part and weak converse, and $E_{1}(n),\,E_{2}(n)\geq\Omega(1/n)$ for the strong converse. By direct substitution of these two “settings” into the results above we can recover the results from [17]. Before proving it, we point out that those results are bounds on the linearithmic capacity (the maximum asymptotic rate), which we can recover as follows in the present notation:
@@
where the inner maximum is over all DI codes with errors of first and second kind bounded as $\lambda_{1}\leq 2^{-nE_{1}(n)}$ and $\lambda_{2}\leq 2^{-nE_{2}(n)}$, and the outer supremum is over functional dependencies of the exponents on $n$ going to 0 sufficiently slowly: $\omega(1/n)\leq E_{1,2}(n)\leq o(1)$.
Theorem 9 (Cf. [17, Thm. 5.7]).
The linearithmic DI capacity $\dot{C}_{\text{DI}}(W)$ for the channel $W:\mathcal{X}\rightarrow\mathcal{Y}$ is bounded in terms of the lower Minkowski dimension of the set $\sqrt{\!\widetilde{\mathcal{X}}}$ as follows (for slowly vanishing error exponents $\lim_{n\rightarrow\infty}E_{i}(n)=0$):
$\frac{1}{4}\underline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\leq\dot{C}_{\text{DI}}(W)\leq\frac{1}{2}\underline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$ (45)
@@
We start with the coding part taking $E_{1}(n)=E_{2}(n)=\frac{\log n}{n}$ in Equation (25), from Corollary 6:
$\begin{split}R(n)&\geq\frac{1-t}{4}\left[\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)-\eta\right]\log\frac{ct^{2}n}{6\log n}\\ &\phantom{\geq.}-H(t,1\!-\!t)-O\!\left(\frac{\log n}{n}\right).\end{split}$ (46)
@@
$\dot{C}_{\text{DI}}(W)\geq\liminf_{n\rightarrow\infty}\frac{1}{\log n}R(n)\geq\frac{1}{4}\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right),$ (47)
@@
Similarly, for the upper bound we need to take $\mathcal{E}_{b}\ni E_{1}(n)=E_{2}(n)\geq\frac{C}{n}$ in Equation (41), the (pessimistic) upper bound for bad sequences of errors. Choosing also $\eta=\eta(n)$ going slowly enough towards 0 for increasing $n$, such that the error exponents, which are upper bounded by the threshold $E_{0}\!\left(\eta(n)\right)$, can also go to 0 slowly, we find
$\begin{split}\dot{C}_{\text{DI}}(W)&\leq\liminf_{n\rightarrow\infty}\frac{1}{2\log n}\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\log\frac{8n}{C}+O\left(\frac{1}{n}\right)\\ &=\frac{1}{2}\underline{d}_{M}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right),\end{split}$ (48)
@@
We refer the reader to [17, Sect. 7.1] for an extensive discussion on the origins, comparison, and interpretation of the optimistic and pessimistic capacities, or else [19] which inspired it. We can recover the general superexponential optimistic capacity bounds in [17] with similar methods as in the previous theorem.
Theorem 10 (Cf. [17, Thm. 5.9]).
The optimistic linearithmic DI capacity $\dot{C}_{\text{DI}}^{\text{opt}}(W)$ of a channel $W:\mathcal{X}\rightarrow\mathcal{Y}$ is bounded in terms of the upper Minkowski dimension of $\sqrt{\!\widetilde{\mathcal{X}}}$ as follows (for slowly vanishing error exponents $\lim_{n\rightarrow\infty}E_{i}(n)=0$):
$\frac{1}{4}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\leq\dot{C}_{\text{DI}}^{\text{opt}}(W)\leq\frac{1}{2}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$ (50)
@@
We follow the same steps as in the previous proof. For the direct part we need to take $\mathcal{E}_{g}\ni E_{1}(n),\,E_{2}(n)\geq\frac{C}{n}$ in the (optimistic) lower bound for good sequences of error exponents [Equation (39)], and similarly choose $t=t(n)$ and $\eta=\eta(n)$ going to 0 simultaneously and sufficiently slowly as $n\rightarrow\infty$. Then,
$\dot{C}_{\text{DI}}^{\text{opt}}(W)$ $\geq\limsup_{n\rightarrow\infty}\frac{1}{4\log n}\left[\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)-\eta(n)\right]\log\frac{ct^{2}n}{6C}$
$=\frac{1}{4}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right).$ (51)
For the converse can we take $E_{1}(n)=E_{2}(n)=\frac{\log n}{n}$ in Equation (35) and $\eta=\eta(n)$ going to zero slowly enough, and find
$\dot{C}_{\text{DI}}^{\text{opt}}(W)$ $\leq\limsup_{n\rightarrow\infty}\frac{1}{2\log n}\left[\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)+\eta(n)\right]\log\frac{8n}{\log n}$
$\phantom{\leq.}+O\left(\frac{1}{n}\right)=\frac{1}{2}\overline{d}_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right),$ (52)
@@
In conclusion, while for exponentially fast vanishing errors the superlinear nature of DI is lost, when imposing a very slow decrease of the errors in block length we can recover the linearithmic capacity bounds in [17] from the rate-reliability bounds derived in Section III. This shows that, despite the slightly superlinear nature of the optimum rates being a general feature of DI, it is also very sensitive to the particular error regimes. The convergence of the rate bounds towards the capacity is also very slow, as shown in Figure 2.
@@ IV-B Zero-dimensional output probability sets
In Corollaries 6 and 8, and in Section III-C, we have demonstrated how to bound the rate $R(n)$ of a code with small error exponents using the lower and upper Minkowski dimensions featured in the main pessimistic and optimistic capacity results of [17]. This could cause some confusion on how to approach channels for which the output probability set (and, equivalently, its square-root) have Minkowski dimension $d_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)=0$, particularly in the regime of small error exponents, where the aforementioned corollaries apply. Indeed, in such a case, the lower bound given in Corollary 6 becomes trivial, and the upper bound only tells us that $R(n)\leq o\left(\log\frac{1}{E(n)}\right)$ for small $E(n)\leq E_{1}(n),\,E_{2}(n)$. However, we can obtain much more meaningful information if we go back to Theorems 4 and 7 (for the lower and upper bounds respectively) and impose the regime of small errors only at an appropriate point.
@@
$B_{x}(y)=xy+(1-x)(1-y)=\begin{cases}x&\text{for }y=1,\\ 1-x&\text{for }y=0.\end{cases}$ (53)
Example 11 (Cf. [17, Subsect. 5.4]).
For a real number $a>1$, let $\mathcal{X}_{a}:=\{0\}\cup\{a^{-k}:0\leq k\in\mathds{N}_{0}\}$ be a set of numbers in the interval [0,1][0,1][ 0 , 1 ]. Then, the maximum rate $R(n)$ of a DI code for $B|_{\mathcal{X}_{a}}$, the Bernoulli channel restricted to inputs from $\mathcal{X}_{a}$, for small and equal error exponents $E(n):=E_{1}(n)=E_{2}(n)>0$, is bounded as follows:
$\log\log\frac{1}{E(n)}-O(1)\leq\max R(n)\leq\log\log\frac{1}{E(n)}+O(1).$ (54)
@@
As the set of all probability distributions with binary output $\mathcal{P}(\{0,1\})$ (with total variation distance) is isometric to the interval [0;1][0;1][ 0 ; 1 ] (with the usual metric on real numbers), the covering (or packing) of the output probability set of a channel with binary output, like $B|_{\mathcal{X}_{a}}$, corresponds to the covering (or packing) of its input, which is $\mathcal{X}_{a}$. It is not hard to see that
@@
To prove the lower bound, we start from Theorem 4, we use the covering/packing relation $\Pi_{\epsilon}(F)\geq\Gamma_{2\epsilon}(F)$, and apply the lower bound in Equation (55), noting that in this example $\sqrt{\mathcal{X}_{a}}=\mathcal{X}_{\sqrt{a}}$. After some basic algebra we get:
$\begin{split}R(n)&\geq(1-t)\log\log_{a}\frac{t\sqrt{c}(\sqrt{a}-1)^{2}}{36\sqrt{6E(n)}}\\ &\phantom{\geq.}-H(t,1-t)-O\left(\frac{\log n}{n}\right).\end{split}$ (56)
To optimise the right-hand side, choose $t=\sqrt[4]{E(n)}$ such that
@@
For the upper bound we start from Theorem 7 and use the right-hand inequality in Equation (55) and that $\sqrt{\mathcal{X}_{a}}=\mathcal{X}_{\sqrt{a}}$:
$\begin{split}R(n)&\leq\log\left[\Gamma_{\frac{1}{2}{\sqrt{1-e^{-E(n)/2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right]\\ &\leq\log\log_{\sqrt{a}}\left(\frac{a}{\sqrt{1-e^{-E(n)/2}}}\right).\end{split}$ (58)
@@
$\begin{split}R(n)&\leq\log\log_{a}\frac{2a^{2}}{E(n)}+O\left(\frac{E(n)}{-\log E(n)}\right)\\ &=\log\log\frac{1}{E(n)}+O(1).\end{split}$ (59)
Considering the bounds above and the arguments in Section IV-A, where we took $E(n)=\omega(1/n)$ for the lower bound and $E(n)=\Omega(1/n)$ for the upper, we conclude that the DI capacity in this case ought be defined at $n\log\log n$ scale (cf. [17, Subsect. 5.4]), and that it evaluates to
@@
The linear capacity results for the DMC from [8, 9] are recovered from these two bounds in the limit of vanishing exponents: $C_{\text{DI}}(W)=\log|W(\mathcal{X})|$.
@@
For the direct part [Equation (61)] we can reuse the proof of Theorem 4 with slight modifications. There, we started by defining a packing on $\sqrt{\!\mathcal{X}}$ of radius $\beta$, and worked after with the discrete set of $|\mathcal{X}_{0}|$ packing elements. Since now we already have a finite set of output probability distributions $\widetilde{\mathcal{X}}=W(\mathcal{X})$, we let
$2\beta:=\min_{x\neq x^{\prime}\in\mathcal{X}}\left|\sqrt{W_{x}}-\sqrt{W_{x^{\prime}}}\right|_{2},$ (63)
[…]
$R(n)\geq(1-t)\log|W(\mathcal{X})|-H(t,1-t)-O\left(\frac{\log n}{n}\right),$ (64)
@@
For the converse [Equation (62)] we start by noticing that Equation (30) translates to a minimum Hamming distance between code words. Namely, letting $\frac{1}{\alpha}:=\displaystyle\max_{x\neq x^{\prime}\in\mathcal{X}}F(W_{x},W_{x^{\prime}})<1$ be the maximum fidelity between different output probability distributions, we have for two distinct code words $u_{j}=x^{n}$ and $u_{k}={x^{\prime}}^{n}$:
$\frac{1}{\alpha^{d_{H}(u_{j},u_{k})}}\leq 4e^{-nE(n)},$ (66)
@@
$R(n)\leq\log|W(\mathcal{X})|-O\left(-E(n)\log E(n)\right).$ (67)
The only thing left to show is that in the asymptotic limit, the rate bounds above yield the linear capacity results from [9]. Repeating the argument in Section IV-A we take $E(n)\geq\Omega(1/n)$ for the upper bound and $E(n)\geq\omega(1/n)$ for the lower, which translates into:
$\begin{split}\max R(n)&\geq\log|W(\mathcal{X})|-O\left(\frac{1}{\sqrt{n}}\log n\right)\\ \max R(n)&\leq\log|W(\mathcal{X})|-O\left(\frac{1}{n}\log n\right).\end{split}$ (68)